OpenStudy (anonymous):

is (6,-6) a solution of 2y-4x=-12

OpenStudy (lgbasallote):

(6,-6)<--6 is the x-coordinate; -6 is the y-coordinate substitute these into 2y - 4x = -12 2(-6) - 4(6) = -12 now see if the left hand side equals the right hand side -12 - 24 = -12 does the left hand side equal the right?

OpenStudy (theviper):

yeah